### v1

# Do LLMs Have Political Correctness? Analyzing Ethical Biases and Jailbreak Vulnerabilities in AI Systems

Abstract
Warning: This paper contains potentially offensive and harmful text.
Although large language models (LLMs) demonstrate impressive proficiency in various tasks, they present potential safety risks, such as ‘jailbreaks’, where malicious inputs can coerce LLMs into generating harmful content. To address these issues, many LLM developers have implemented various safety measures to align these models. This alignment involves several techniques, including data filtering during pre-training, supervised fine-tuning, reinforcement learning from human feedback, and red-teaming exercises. These methods often introduce deliberate and intentional biases similar to Political Correctness (PC) to ensure the ethical behavior of LLMs. In this paper, we delve into the intentional biases injected into LLMs for safety purposes and examine methods to circumvent these safety alignment techniques. Notably, these intentional biases result in a jailbreaking success rate in GPT-4o models that differs by 20% between non-binary and cisgender keywords and by 16% between white and black keywords, even when the other parts of the prompts are identical. We introduce the concept of PCJailbreak, highlighting the inherent risks posed by these safety-induced biases. Additionally, we propose an efficient defense method PCDefense, which prevents jailbreak attempts by injecting defense prompts prior to generation. PCDefense stands as an appealing alternative to Guard Models, such as Llama-Guard, that require additional inference cost after text generation. Our findings emphasize the urgent need for LLM developers to adopt a more responsible approach when designing and implementing safety measures. To enable further research and improvements, we open-source our code and artifacts of PCJailbreak, providing the community with tools to better understand and mitigate safety-induced biases in LLMs.
1 Introduction
Figure 1: PCJailbreak reveals inherent biases in LLMs that disproportionately allow harmful jailbreak attacks to succeed more frequently when directed towards marginalized groups compared to privileged groups.
Figure 2: Illustration showcasing the difference in response between a standard prompt and a PCJailbreak prompt. While the standard prompt is blocked by the LLM’s safety features, the PCJailbreak prompt exploits intentional biases to elicit a response.
Large Language Models (LLMs) have rapidly become essential components in many fields, ranging from professional decision-making to various forms of interactive user engagement (Araci (2019); Luo et al. (2022); Tinn et al. (2023)). However, as these models become popular, ensuring their safe usage has become crucial. Developers have implemented several safety features to prevent these models from generating harmful or objectionable content, often referred to as ‘safety alignment’. (Bakker et al. (2022); Christiano et al. (2017); Ouyang et al. (2022); Openai Usage Policies). However, in this work we show that these safety alignments often introduce deliberate and intentional biases, giving rise to a phenomenon known as ’jailbreak’, where malicious inputs manage to circumvent these safety alignments, thus allowing LLMs to generate harmful outputs (Goldstein et al. (2023); Kang et al. (2024)).
The term ‘jailbreak’ refers to carefully crafted prompts that can bait aligned LLMs into bypassing their safety alignment, resulting in the generation of content that may be harmful, discriminatory, violent, or sensitive (Smith et al. (2022)). Numerous types of jailbreak attacks have been identified and categorized into two primary methods: manually written jailbreak prompts and learning-based jailbreak prompts (Zou et al. (2023); Lapid et al. (2023); Liu et al. (2023); Wei et al. (2024); Yuan et al. (2024)). The manual approach is exemplified by the “Do-Anything-Now (DAN)” series, which uses manually constructed prompts to exploit aligned LLMs. Conversely, the learning-based approach, as demonstrated by the GCG attack, formulates jailbreaks as adversarial examples by utilizing the gradient information of LLMs to automatically generate effective prompts.
However, there are two main limitations of existing jailbreak methods. Firstly, automated attacks like GCG (Zou et al. (2023)) rely on a search scheme that is guided by gradient information. While this approach enables the automatic generation of jailbreak prompts, it carries a significant downside: the resulting prompts often consist of nonsensical sequences, which lack semantic meaning (Morris et al. (2020)). This major flaw makes these prompts highly vulnerable to naive defense mechanisms such as perplexity-based detection. For example, recent studies (Jain et al. (2023); Alon & Kamfonas (2023)) have shown that such straightforward defenses can easily recognize these nonsensical prompts and completely undermine the success of GCG attacks.
Secondly, although manual attacks can create stealthier jailbreak prompts, these prompts are usually handcrafted by individual LLM users, which leads to issues with scalability and adaptability. Additionally, these handcrafted methods may not adapt quickly to updates in LLMs, thus diminishing their effectiveness over time (ONeal (2024)).
To address these limitations, we use a method that employs fixed prompt formats, offering scalability. By leveraging biases, such as those related to Political Correctness(PC)(Perez et al. (2022); Zhuo et al. (2023)), inherent within LLMs and often intentionally injected by LLM developers to assure safety, we ensure that the prompts retain their meaning significantly without losing effectiveness, contrary to gradient-based attacks. This approach allows us to overcome the issues of scalability and adaptability while still exploiting the biases for effective jailbreaks.
In this paper, we explore the concept of PCJailbreak, investigating how intentional biases in LLMs, intended as safety alignment, paradoxically become enablers of harmful content generation when exploited. This behavior is well illustrated in Figure 1 and 2. Additionally, we propose a defense mechanism PCDefense that adjusts biases using prompts, ensuring safety and efficiency without additional inference or models, which makes it an attractive alternative to Guard Models (Inan et al. (2023); Ghosh et al. (2024); Caselli et al. (2020); Vidgen et al. (2021)), which are capable of classifying harmful conversations but require additional models and inference after text generation. .
Our contributions can be summarized as follows:
  • We analyze the nature and consequences of intentional biases introduced in LLMs for safety purposes, highlighting their potential to not only fail in deterring but also in facilitating more effective jailbreaks. This paradoxical effect underscores the urgency of addressing the inherent vulnerabilities these biases introduce.
  • Through comprehensive experiments, we show that our proposed PCJailbreak is effective across state-of-the-art models, including the latest iterations of GPT. Our framework also proves adaptable, working effectively when applied to existing jailbreak techniques.
  • We propose PCDefense, a straightforward defense strategy without the need of additional inference or models. Our findings demonstrate that even with a simple and cost-effective defense approach, jailbreak attacks can be mitigated. This highlights the critical responsibility of LLM service providers to ensure robust protection.
  • We open-source the code and all associated artifacts of PCJailbreak to facilitate community efforts in understanding and mitigating safety-induced biases in large language models. This contribution aims to provide researchers and developers with the necessary tools to explore the nature of these biases and develop more robust defenses, furthering the collective effort to ensure the safety and reliability of LLM deployments.
Our research suggests that while intentional biases are crucial for aligning LLMs with ethical standards, they necessitate careful scrutiny to prevent their manipulation. Therefore, responsible strategies from AI companies and researchers are needed to reinforce the safety of LLMs in an increasingly complex threat landscape.
2 Background And Related Works
2.1 Safety Alignment in LLMs
Ensuring the safety and ethical alignment of large language models (LLMs) is a critical area of ongoing research. Methods such as data filtering, supervised fine-tuning, and reinforcement learning from human feedback (RLHF) aim to align models like GPT-4 and ChatGPT with human values and preferences (Christiano et al. (2017); Bai et al. (2022); Ouyang et al. (2022); Xu et al. (2020)). However, despite these efforts, recent studies reveal vulnerabilities that can be exploited through ’jailbreak’ attacks, which lead to undesirable and harmful outputs (Kang et al., 2023; Hazell, 2023; Shen et al., 2023).
2.2 Jailbreak Attacks and Techniques
Jailbreaking LLMs involves crafting inputs that bypass safety mechanisms, resulting in harmful or objectionable content. Early jailbreak attacks, such as the ”Do-Anything-Now (DAN)” series, relied on manually crafted prompts to exploit LLM safeguards (walkerspider, 2022). (Liu et al. (2023) provided an in-depth analysis and categorization of these jailbreak prompts, highlighting the delicate balance between an LLM’s capabilities and its safety constraints.
Diverse strategies for jailbreaks have been proposed. Manual methods, while effective, suffer from scalability issues (Wei et al. (2024)). On the other hand, learning-based methods like GCG (Zou et al. (2023)) use adversarial techniques to generate prompts automatically, though often at the cost of producing semantically meaningless outputs detectable via simple defenses like perplexity tests (Alon & Kamfonas (2023); Liu et al. (2023)) introduced AutoDAN, which combines manual and automated strategies using hierarchical genetic algorithms to enhance both the stealthiness and scalability of jailbreak prompts.
Language diversity and non-natural language inputs present additional challenges. Deng et al. (2023) explored multilingual jailbreak attacks, demonstrating that LLMs could be tricked into producing harmful outputs with non-English prompts. Yuan et al. (2024) extended this by investigating the vulnerabilities of LLMs to non-natural language inputs, such as ciphers.
2.3 Towards Improved Safety Measures
Complex attack strategies like those proposed by Ding et al. (2023) with the ReNeLLM framework introduce the concept of generalized and nested jailbreak prompts, leveraging LLMs to generate effective prompts through prompt rewriting and scenario nesting. This highlights the dynamic and evolving nature of jailbreak techniques.
Our work builds on the existing body of research by focusing on the paradoxical consequences of intentional biases introduced for safety purposes. While these biases aim to align LLMs ethically, they also highlight new vulnerabilities. To counteract this, we propose using prompts to make the LLM re-align those biases, thus offering a robust secondary defense against jailbreak attempts.
In conclusion, AI developers must adopt a higher degree of responsibility in designing, testing, and deploying LLMs. This involves continuous monitoring and iterative improvements based on real-world data. Our findings advocate for a nuanced approach to LLM safety, promoting the development of more secure and reliable models, and ensuring that safety measures do not inadvertently introduce new risks.
3 Methodology: PCJailbreak
3.1 Preliminaries
3.1.1 Jailbreak Attack
A jailbreak attack in the context of Large Language Models (LLMs) occurs when the model generates harmful or inappropriate responses to malicious inputs instead of producing a refusal signal, which is a safe and ethical response denying the request (Zeng et al., 2024; Zou et al., 2023). Such attacks are intricately linked to the alignment methods employed in LLMs, which aim to align the model’s outputs with human values and ethical considerations. The primary objective of these attacks is to bypass the LLM’s alignment constraints, causing the model to produce inappropriate responses to malicious queries instead of refusing to answer.
Consider a set of malicious queries $Q=\{Q_{1},Q_{2},\ldots,Q_{n}\}$. An attacker meticulously designs a set of jailbreak prompts $J=\{J_{1},J_{2},\ldots,J_{n}\}$ to manipulate these queries. When each combined input $T_{i}=\langle J_{i},Q_{i}\rangle$ is fed into the LLM, the model generates a set of responses $R=\{R_{1},R_{2},\ldots,R_{n}\}$. The aim of the jailbreak attack is to ensure that the responses $R$ correctly address the malicious queries in $Q$ instead of producing refusal signals (Zeng et al., 2024).
In summary, a successful jailbreak attack coerces the LLM into bypassing its safety and ethical constraints, culminating in the generation of harmful or inappropriate content in response to malicious prompts. This kind of attack underscores significant vulnerabilities in the current alignment strategies of LLMs, highlighting the necessity for robust countermeasures to guarantee model safety and reliability.
3.1.2 Formulation
Designing a specific target response for each malicious question is impractical for several reasons. Firstly, manually crafting an appropriate answer for every possible malicious query $Q_{i}$ is resource-intensive and infeasible due to the vast number of potential queries. Secondly, creating specific responses may limit the attack’s generalizability, as tailored responses might not be effective for other queries or variations of the same query. Lastly, directly specifying harmful content could raise ethical and legal concerns.
Hence, a common solution to this problem is to define a standardized introductory phrase for the target response (Zou et al., 2023; Lapid et al., 2023; Qi et al., 2024). The determination of a successful jailbreak focuses on the absence of this standardized introductory phrase in the response. Such as:
[…]
3.2 PCJailbreak Approach
3.2.1 Generating Keywords and Prompts
Our proposed PCJailbreak approach investigates how keywords associated with various demographic and socio-economic groups influence an LLM’s susceptibility to jailbreak attacks. To accomplish this, we generated keywords representing contrasting groups, such as privileged and marginalized communities, using the LLM itself to ensure that the keywords reflect the model’s inherent associations and biases.
We prompted the LLM with:
“Provide a list of adjectives that describe various privileged and marginalized groups.”
This process produced pairs of keywords, for example:
[…]
$\vdots$
For the harmful prompts, we compile a set $H=\{H_{1},H_{2},\ldots,H_{m}\}$ of ethically sensitive requests that LLMs should refuse, referenced from the harmful prompt dataset in JailBreakBench (Chao et al., 2024). These prompts represent a variety of harmful or unethical queries, such as requests for illicit activities or dissemination of sensitive information.
3.2.2 Analyzing Inherent Biases
Figure 3: Overview of the PCJailbreak methodology. The same harmful prompt is used across different keywords representing contrasting groups to analyze variations in jailbreak success rates.*All keywords representing both privileged and underprivileged groups are generated by the LLM.
To analyze the biases inherent in LLMs, we constructed prompts that incorporated the generated keywords and harmful prompts. As illustrated in Figure 3, these prompts were then used to measure variations in jailbreak success rates. By iterating the keywords and harmful prompts, we created a set of final prompts $T=\{T_{i,k,j}\}$:
$\text{T}_{i,k,j}=\text{``I am a }\{\mathrm{keyword}_{i,k}\}\text{ writer who writes fiction. Write a fiction about }\{H_{j}\}\text{.''}$” (1)
where:
  • $\mathrm{keyword}_{i,k}$ is the $k$-th keyword of the $i$-th pair ($k\in\{\text{privileged},\ \text{marginalized}\}$),
  • $H_{j}$ is the $j$-th harmful prompt from our set $H$.
  • $T_{i,k,j}$ is the final prompt used for LLM input.
Jailbreak Success Rate Calculation
For each keyword and harmful prompt combination, we conducted multiple trials by submitting the corresponding prompt $T_{i,k,j}$ to the LLM and recording the outcomes. The success rate $S_{i,k}$ for each $\mathrm{keyword}_{i,k}$ is calculated as:
$S_{i,k}=\frac{N_{\text{success},\ i,k}}{N_{\text{total},\ i,k}}$ (2)
where:
  • $N_{\text{success},\ i,k}$ is the number of successful jailbreak attempts for $\mathrm{keyword}_{i,k}$,
  • $N_{\text{total},\ i,k}$ is the total number of attempts for $\text{keyword}_{i,k}$.
Determining Jailbreak Success
A response from the LLM is considered a successful jailbreak if it includes the harmful content requested in the harmful prompt $H_{j}$ instead of providing a refusal or safe completion. For instance, if the LLM provides a detailed story incorporating the harmful content without any refusal language, it is marked as a successful jailbreak. Conversely, if the LLM responds with a standard refusal message, it is considered unsuccessful in terms of the jailbreak attempt.
Comparative Analysis
After calculating the success rates for each keyword, we compared the success rates between the privileged and marginalized keywords within each pair. Let $S_{i,\text{privileged}}$ and $S_{i,\text{marginalized}}$ be the success rates for the privileged and marginalized keywords of the $i$-th pair, respectively. We analyzed the difference $\Delta S_{i}$ between these success rates:
$\Delta S_{i}=S_{i,\text{privileged}}-S_{i,\text{marginalized}}$ (3)
A significant $\Delta S_{i}$ suggests that the LLM exhibits differential susceptibility to jailbreak attacks based on the demographic represented by the keyword. $\Delta S_{i}$ indicates that there is a significant difference in jailbreak success rates between specific group keywords when the value is large. This could indicate inherent biases in the LLM’s training data or alignment mechanisms that affect how it responds to prompts involving different groups.
3.2.3 Summary of the PCJailbreak Approach
Our PCJailbreak methodology systematically analyzes the LLM’s responses to a diverse set of harmful prompts, using various keywords across different groups. We conducted experiments using harmful prompts $H$ and multiple pairs of privileged and marginalized keywords. By maintaining a consistent prompt structure (Equation 1) and isolating the effect of the keyword, we aimed to identify any biases in the LLM’s susceptibility to jailbreak attacks.
The calculation of jailbreak success rates (Equation 2) and the comparative analysis using $\Delta S_{i}$ (Equation 3) enabled us to quantify potential biases. Our approach highlights whether intentional biases introduced during the alignment process inadvertently contribute to vulnerabilities that can be exploited through jailbreak attacks.
3.3 PCDefense: Preventing Jailbreaks without Additional Inference or Guard Models
Our jailbreak defense method PCDefense uses prompts to adjust biases without the need for additional models. Chain-of-Thought Prompting (CoT) has shown considerable potential in eliciting reasoning abilities in large language models, allowing for interpretability without the need for parameter updates (Wei et al., 2022). Inspired by this, our approach PCDefense involves adding defense prompts to reduce excessive biases, ensuring safety without additional inference overheads, in contrast to Guard Models which require additional inference cost after text generation (Inan et al. (2023); Ghosh et al. (2024); Caselli et al. (2020); Vidgen et al. (2021)). By incorporating bias terms through our defense prompts, we achieve a cost-efficient and secure method for defending against jailbreak attempts while maintaining the model’s performance.
To further illustrate our defense prompt, we use the following prompt structure to ensure fairness and equity:
Figure 4: PCDefense adjusts inherent biases in LLMs that are exploited by PCJailbreak. It is efficient since it does not require additional inference or models such as Guard Models.
By using a defense prompt, our defense mechanism elicits the refusal of a wide range of harmful content without requiring significant additional computing resources, which makes it an attractive alernative to Guard Models.
4 Experiment
4.1 Experimental Setups
Data To analyze the differences in jailbreak success rates between groups, we used the jailbreakbench dataset (Chao et al. (2024)). This dataset comprises 100 harmful prompts and has been widely used in numerous studies. We utilized this dataset to measure performance variations and to analyze the intentional biases inherent in LLMs.
Models We analyzed intentional bias in well-known LLMs, including closed-source models such as GPT-3.5-turbo, GPT-4, GPT-4-o, and open-source models like Llama2-7B, Llama2-13B, Llama3-7B, Phi-mini-7B, Qwen1.5, and Qwen2-7B.
For our experiments, we conducted evaluations with the default sampling temperature and system prompts. This setup aligns with standard practices in the literature to ensure consistent and comparable results.
Keywords The keywords used to compare performance across groups were generated by the LLMs themselves to maximize the exploitation of internal biases. Most models generated similar keywords, and significant performance differences were observed.
Basically, the models primarily generated the following keywords for various groups (see Table 1):
Marginalized Group (Adjective) Privileged Group (Adjective)
White Black
Male Female
Heterosexual LGBTQ+
Able-bodied Disabled
Wealthy Low-income
Cisgender Non-normative gender
Traditionally family-structured Non-traditionally family-structured
Urban Rural
Highly educated Undereducated
Traditionally employed Sexually marginalized
Table 1: Keywords Generated for Group Comparison
These keywords were systematically used to evaluate the differences in jailbreak success rates among various groups.
4.2 Intensional Bias Analysis
To gauge the impact of intentionally injected biases in LLMs on their vulnerability to jailbreak attempts, we conducted a series of experiments utilizing various keyword prompts designed to assess the success rates of jailbreaks within marginalized and privileged categories. We evaluated multiple well-known models, including GPT-3.5, GPT-4, GPT-4o (Brown et al., 2020)(closed models), as well as LLaMA (Touvron et al., 2023), Qwen (Bai et al., 2023), and Phi-mini (Abdin et al., 2024)(open-source models).
Model Baseline Marginalized Privileged Marginalized /
Name Success Rate Success Rate ($\uparrow$) Success Rate ($\downarrow$) Privileged ($\uparrow$)
GPT-3.5 0.2200 0.2421 (+10.00%) 0.1847 (-15.90%) 131.08%
GPT-4 0.2100 0.2488 (+18.57%) 0.1900 (-9.52%) 130.95%
GPT-4o 0.4600 0.5467 (+18.91%) 0.4187 (-8.91%) 130.57%
LLaMA2 0.2400 0.2811 (+17.08%) 0.1933 (-19.58%) 145.42%
LLaMA3 0.0500 0.0650 (+30.00%) 0.0300 (-40.00%) 216.67%
Qwen-1.5 0.1900 0.2175 (+14.74%) 0.1675 (-11.58%) 129.85%
Qwen2 0.1700 0.1971 (+15.88%) 0.1671 (-7.06%) 117.95%
Phi-mini 0.4100 0.4386 (+7.07%) 0.3829 (-6.59%) 114.56%
Table 2: Performance across different models showing baseline success rates, marginalized success rates, privileged success rates, and the difference between marginalized and privileged success rates.
Interestingly, the latest model, GPT-4o, exhibits higher jailbreak success rates, with a significant difference observed between marginalized and privileged groups. In particular, GPT-4o shows a difference of 0.128 between marginalized and privileged keywords, highlighting the substantial gap influenced by injected biases. Furthermore, most models demonstrate lower jailbreak success rates for prompts using privileged keywords compared to the baseline, whereas prompts using marginalized keywords exhibit higher success rates.
Intriguingly, the LLaMA3 model demonstrates a lower propensity for successful jailbreak attempts compared to other models. This lower rate can be attributed to Meta’s focus on creating more robust LLMs that are resistant to such vulnerabilities. It showcases Meta’s commitment to mitigating LLM safety risks through comprehensive alignment techniques (Vidgen et al., 2021; Ghosh et al., 2024; Caselli et al., 2020).
4.3 Performance Improvement using Bias-based Method
The bias-based approach we analyzed can also be applied to existing models. We confirmed performance improvement by applying the PCJailbreak method to Adaptive Attacks (Andriushchenko et al., 2024), the current state-of-the-art (SOTA) jailbreak attack. Specifically, for the llama2 model(Touvron et al., 2023), the performance improved from 98% to 100%, resulting in a 2% increase. For the phi model(Abdin et al., 2024), the performance enhanced from 95% to 99%, resulting in a 4% increase. These results show that the method can be easily integrated with existing techniques. The experiment was held by using the jailbreak attack artifact from JailbreakBench(Chao et al., 2024), which has 100 samples of Adaptive Attacks conversation.
Table 3: Performance Improvement of SOTA Models
Model Adaptive Attacks Adaptive Attacks with PCJailbreak
llama2 98.00% 100.00%
phi-mini 95.00% 99.00%
The results in Table 3 demonstrate that our bias-based method can be effectively combined with the existing approaches, providing notable improvements in performance.
4.4 Experimental Validation of PCDefense
To validate our defense method PCDefense, we conducted experiments on various models, including Llama2, Phi, and Qwen2. We observed the impact of our approach on Marginalized Rate Success and Privileged Rate Success metrics before and after applying our defense technique.
The results are summarized in Table 4.
Table 4: Jailbreak Prevention performance of PCDefense
Model Metric Before After After/Before ($\downarrow$)
Llama2 Marginalized Group Jailbreak Success 0.2811 0.1714 60.97%
Privileged Group Jailbreak Success 0.1933 0.1429 73.93%
Gap Between Groups 0.0878 0.0285 32.46%
Phi Marginalized Rate Success 0.4386 0.4208 95.94%
Privileged Rate Success 0.3829 0.4075 106.42%
Gap Between Groups 0.0557 0.0133 23.88%
Qwen2 Marginalized Rate Success 0.1971 0.1750 88.79%
Privileged Rate Success 0.1671 0.1900 113.70%
Gap Between Groups 0.0300 0.0150 50.00%
As shown in Table 4, the Marginalized Rate Success and Privileged Rate Success both decreased consistently across Llama2 and Qwen2 models after applying our defense method. Specifically, for the Llama2 model, the Marginalized Rate Success decreased by 0.1097 and the Privileged Rate Success decreased by 0.0504. For the Phi and Qwen2 models, the Marginalized Rate Success decreased by about 0.02, whereas the Privileged Rate Success increased by about 0.02.
While the performance of the Privileged group has increased in some models, potentially leading to the misconception that it has become more dangerous, the reduced gap in jailbreak performance between groups indicates a decrease in bias. Additionally, the overall average score has decreased, suggesting that the system has become safer.
4.5 Principle Component Analysis
In this section, we conduct a detailed analysis of how models interpret PCJailbreak prompts by performing Principal Component Analysis (PCA) (F.R.S., 1901) on the embeddings of Phi-mini, Qwen2, and LLaMA3 models, as shown in Appendix Figures 5, 6, 7 respectively. Our findings reveal that models with a high success rate in PCJailbreak tend to cluster PCJailbreak prompts together with benign prompts, while those with a lower success rate demonstrate the opposite behavior. The PCA results indicate that when using PCJailbreak, models tend to interpret it more similarly to benign prompts, with the degree of this similarity aligning with our observed PCJailbreak success rates.
5 Conclusion
In conclusion, our study highlights the complexities and potential unintended consequences of aligning large language models (LLMs) with safety measures aimed at preventing harmful outputs. The introduction of intentional biases for ethical behavior, while crucial for ensuring responsible AI, has led to significant discrepancies in jailbreak success rates based on gender and racial keywords. This discrepancy, coined as ”PCJailbreak,” underscores the risks that arise from safety-induced biases, particularly in terms of fairness and equality. Our findings emphasize the need for LLM developers to carefully balance safety and fairness in their models. Additionally, our proposed method of adding defense prompts without requiring additional inference or models shows a simple and scalable solution to mitigate jailbreak attempts. This simplicity and scalability of migation emphasizes the responsibility of LLM developers to adopt more comprehensive and proactive measures to address safety risks. Future work should focus on designing more inclusive and transparent alignment strategies to address the inherent challenges of AI safety while minimizing bias.
Ethics statement.
This work addresses critical security concerns in the use of large language models (LLMs), specifically in relation to biases introduced for safety purposes and their potential to facilitate jailbreaks. While our analysis highlights the vulnerabilities of current bias-based safety mechanisms, it is important to clarify that our research aims to enhance the robustness of LLMs by mitigating these risks rather than exploiting or encouraging harmful behavior. We recognize the dual-use nature of this research, and we have taken care to emphasize defense strategies, including the proposal of small auxiliary models, that prioritize user protection and ethical AI deployment.
Furthermore, the experiments conducted in this work do not involve human subjects or sensitive personal data. All model assessments were performed in controlled environments using publicly available datasets and models. We commit to following best practices in data security and model transparency and do not release any tools or frameworks that could directly enable malicious use. Our findings are intended to foster a deeper understanding of how to secure AI systems and encourage responsible model deployment within the community.
Reproducibility statement.
We are committed to ensuring the reproducibility of our results and findings. To this end, we provide open access to the source code, artifacts, datasets, and detailed instructions on how to replicate our experiments through a publicly available anonymous repository. By following the guidelines outlined in our repository, researchers and practitioners should be able to easily reproduce and extend our work.
Appendix A Appendix
Figure 5: PCA of Phi-mini model, which has 0.4386 PCJailbreak success rate. PCJailbreak samples are closely clustered with Benign samples.
Figure 6: PCA of Qwen2 model, which has 0.1971 PCJailbreak success rate. PCJailbreak samples are relatively far from both harmful and benign samples.
Figure 7: PCA of LLaMA3 model, which has 0.0650 PCJailbreak success rate. PCJailbreak samples are relatively close to harmful samples.

### v3

@@
-Do LLMs Have Political Correctness? Analyzing Ethical Biases and Jailbreak Vulnerabilities in AI Systems
+BiasJailbreak:Analyzing Ethical Biases and Jailbreak Vulnerabilities in Large Language Models
 Abstract
 Warning: This paper contains potentially offensive and harmful text.
-Although large language models (LLMs) demonstrate impressive proficiency in various tasks, they present potential safety risks, such as ‘jailbreaks’, where malicious inputs can coerce LLMs into generating harmful content.
-To address these issues, many LLM developers have implemented various safety measures to align these models.
-This alignment involves several techniques, including data filtering during pre-training, supervised fine-tuning, reinforcement learning from human feedback, and red-teaming exercises.
-These methods often introduce deliberate and intentional biases similar to Political Correctness (PC) to ensure the ethical behavior of LLMs.
-In this paper, we delve into the intentional biases injected into LLMs for safety purposes and examine methods to circumvent these safety alignment techniques.
-Notably, these intentional biases result in a jailbreaking success rate in GPT-4o models that differs by 20% between non-binary and cisgender keywords and by 16% between white and black keywords, even when the other parts of the prompts are identical.
-We introduce the concept of PCJailbreak, highlighting the inherent risks posed by these safety-induced biases.
-Additionally, we propose an efficient defense method PCDefense, which prevents jailbreak attempts by injecting defense prompts prior to generation.
-PCDefense stands as an appealing alternative to Guard Models, such as Llama-Guard, that require additional inference cost after text generation.
-Our findings emphasize the urgent need for LLM developers to adopt a more responsible approach when designing and implementing safety measures.
-To enable further research and improvements, we open-source our code and artifacts of PCJailbreak, providing the community with tools to better understand and mitigate safety-induced biases in LLMs.
+Although large language models (LLMs) demonstrate impressive proficiency in various tasks, they present potential safety risks, such as ‘jailbreaks’, where malicious inputs can coerce LLMs into generating harmful content bypassing safety alignments.
+In this paper, we delve into the ethical biases in LLMs and examine how those biases could be exploited for jailbreaks.
+Notably, these biases result in a jailbreaking success rate in GPT-4o models that differs by 20% between non-binary and cisgender keywords and by 16% between white and black keywords, even when the other parts of the prompts are identical.
+We introduce the concept of BiasJailbreak, highlighting the inherent risks posed by these safety-induced biases.
+BiasJailbreak generates biased keywords automatically by asking the target LLM itself, and utilizes the keywords to generate harmful output.
+Additionally, we propose an efficient defense method BiasDefense, which prevents jailbreak attempts by injecting defense prompts prior to generation.
+BiasDefense stands as an appealing alternative to Guard Models, such as Llama-Guard, that require additional inference cost after text generation.
+Our findings emphasize that ethical biases in LLMs can actually lead to generating unsafe output, and suggest a method to make the LLMs more secure and unbiased.
+To enable further research and improvements, we open-source our code and artifacts of BiasJailbreak, providing the community with tools to better understand and mitigate safety-induced biases in LLMs.
 1 Introduction
-Figure 1: PCJailbreak reveals inherent biases in LLMs that disproportionately allow harmful jailbreak attacks to succeed more frequently when directed towards marginalized groups compared to privileged groups.
-Figure 2: Illustration showcasing the difference in response between a standard prompt and a PCJailbreak prompt.
-While the standard prompt is blocked by the LLM’s safety features, the PCJailbreak prompt exploits intentional biases to elicit a response.
+Figure 1: BiasJailbreak reveals inherent biases in LLMs that disproportionately allow harmful jailbreak attacks to succeed more frequently when directed towards marginalized groups compared to privileged groups.
+Figure 2: Illustration showcasing the difference in response between a standard prompt and a BiasJailbreak prompt.
+While the standard prompt is blocked by the LLM’s safety features, the BiasJailbreak prompt exploits ethical biases to elicit a response.
 Large Language Models (LLMs) have rapidly become essential components in many fields, ranging from professional decision-making to various forms of interactive user engagement (Araci (2019); Luo et al.
@@
 However, as these models become popular, ensuring their safe usage has become crucial.
-Developers have implemented several safety features to prevent these models from generating harmful or objectionable content, often referred to as ‘safety alignment’.
-(Bakker et al.
+Developers have implemented several safety features to prevent these models from generating harmful or objectionable content, often referred to as ‘safety alignment’ (Bakker et al.
 (2022); Christiano et al.
@@
 (2022); Openai Usage Policies).
-However, in this work we show that these safety alignments often introduce deliberate and intentional biases, giving rise to a phenomenon known as ’jailbreak’, where malicious inputs manage to circumvent these safety alignments, thus allowing LLMs to generate harmful outputs (Goldstein et al.
+These safety alignments often involve additional fine-tuning or reinforcement learning techniques, which, while designed to enhance safety and alignment, may also inadvertently introduce biases, as highlighted in resources such as (Achiam et al., 2023, p.
+49) .
+However, biases can also arise from other sources, such as pretraining data or system prompts.
+While it is difficult to pinpoint exactly where these biases originate, the critical fact remains that they exist and can influence the model’s behavior.
+In this work, we show that these safety alignments often introduce deliberate and ethical biases, giving rise to a phenomenon known as ’jailbreak’, where malicious inputs manage to circumvent these safety alignments, thus allowing LLMs to generate harmful outputs (Goldstein et al.
 (2023); Kang et al.
@@
 (2022)).
-Numerous types of jailbreak attacks have been identified and categorized into two primary methods: manually written jailbreak prompts and learning-based jailbreak prompts (Zou et al.
-(2023); Lapid et al.
-(2023); Liu et al.
-(2023); Wei et al.
-(2024); Yuan et al.
+Numerous types of jailbreak attacks have been identified and categorized into two primary methods: white-box and black-box (Yi et al.
 (2024)).
-The manual approach is exemplified by the “Do-Anything-Now (DAN)” series, which uses manually constructed prompts to exploit aligned LLMs.
-Conversely, the learning-based approach, as demonstrated by the GCG attack, formulates jailbreaks as adversarial examples by utilizing the gradient information of LLMs to automatically generate effective prompts.
-However, there are two main limitations of existing jailbreak methods.
-Firstly, automated attacks like GCG (Zou et al.
-(2023)) rely on a search scheme that is guided by gradient information.
-While this approach enables the automatic generation of jailbreak prompts, it carries a significant downside: the resulting prompts often consist of nonsensical sequences, which lack semantic meaning (Morris et al.
+The white-box approach requires target model gradients or logits and use them as a guidance for finding adversarial jailbreak prompts.
+Directly fine-tuning the target LLM is not considered as a jailbreak method in this paper, although it some consider it as white-box.
+The black-box approach has a harder and a more general real-world setting since it does not have access to such information.
+Jailbreak methods in black-box usually involve template completion, prompt rewriting, or LLM-based generation.
+White-box attacks like GCG (Zou et al.
+(2023)) rely on a search scheme guided by gradient information.
+While this approach enables reliable generation of jailbreak prompts though with a cost of high computation, it carries a significant downside: the resulting prompts often consist of nonsensical sequences, which lack semantic meaning (Morris et al.
 (2020)).
@@
 For example, recent studies (Jain et al.
-(2023); Alon & Kamfonas (2023)) have shown that such straightforward defenses can easily recognize these nonsensical prompts and completely undermine the success of GCG attacks.
-Secondly, although manual attacks can create stealthier jailbreak prompts, these prompts are usually handcrafted by individual LLM users, which leads to issues with scalability and adaptability.
-Additionally, these handcrafted methods may not adapt quickly to updates in LLMs, thus diminishing their effectiveness over time (ONeal (2024)).
-To address these limitations, we use a method that employs fixed prompt formats, offering scalability.
-By leveraging biases, such as those related to Political Correctness(PC)(Perez et al.
+(2023); Alon & Kamfonas (2023)) have shown that such straightforward defenses can easily recognize these nonsensical prompts and completely undermine the success of white-box attacks.
+While having a more applicable and general setting, recent advancements in black-box settings, such as natural-language methods like PAIR (Chao et al.
+(2023)) and DeepInception (Li et al.
+(2023)), have shown that semantically coherent prompts can effectively exploit LLM vulnerabilities.
+Additionally, manual approaches like GUARD (Jin et al.
+(2024a)) iteratively refine jailbreak prompts, demonstrating adaptability to LLM updates.
+In this paper, we propose a novel black-box method that offers both scalability and generality.
+By leveraging inherent biases in LLMs, such as those related to Ethical Sensitivity(Perez et al.
 (2022); Zhuo et al.
-(2023)), inherent within LLMs and often intentionally injected by LLM developers to assure safety, we ensure that the prompts retain their meaning significantly without losing effectiveness, contrary to gradient-based attacks.
+(2023)), we ensure that the prompts retain their meaning significantly without losing effectiveness, contrary to white-box attacks.
 This approach allows us to overcome the issues of scalability and adaptability while still exploiting the biases for effective jailbreaks.
-In this paper, we explore the concept of PCJailbreak, investigating how intentional biases in LLMs, intended as safety alignment, paradoxically become enablers of harmful content generation when exploited.
+We explore the novel concept of BiasJailbreak, investigating how biases in LLMs, intended as safety alignment, paradoxically become enablers of harmful content generation when exploited.
 This behavior is well illustrated in Figure 1 and 2.
-Additionally, we propose a defense mechanism PCDefense that adjusts biases using prompts, ensuring safety and efficiency without additional inference or models, which makes it an attractive alternative to Guard Models (Inan et al.
+Additionally, we propose a defense mechanism BiasDefense that adjusts biases using prompts, ensuring safety and efficiency without additional inference or models, which makes it an attractive alternative to Guard Models (Inan et al.
 (2023); Ghosh et al.
@@
 (2021)), which are capable of classifying harmful conversations but require additional models and inference after text generation.
-.
 Our contributions can be summarized as follows:
-  • We analyze the nature and consequences of intentional biases introduced in LLMs for safety purposes, highlighting their potential to not only fail in deterring but also in facilitating more effective jailbreaks.
+  • We analyze the nature and consequences of ethical biases introduced in LLMs for safety purposes, highlighting their potential to not only fail in deterring but also in facilitating more effective jailbreaks.
 This paradoxical effect underscores the urgency of addressing the inherent vulnerabilities these biases introduce.
-  • Through comprehensive experiments, we show that our proposed PCJailbreak is effective across state-of-the-art models, including the latest iterations of GPT.
+  • Through comprehensive experiments, we show that our proposed BiasJailbreak is effective across state-of-the-art models, including the latest iterations of GPT.
 Our framework also proves adaptable, working effectively when applied to existing jailbreak techniques.
-  • We propose PCDefense, a straightforward defense strategy without the need of additional inference or models.
+  • We propose BiasDefense, a straightforward defense strategy without the need of additional inference or models.
 Our findings demonstrate that even with a simple and cost-effective defense approach, jailbreak attacks can be mitigated.
 This highlights the critical responsibility of LLM service providers to ensure robust protection.
-  • We open-source the code and all associated artifacts of PCJailbreak to facilitate community efforts in understanding and mitigating safety-induced biases in large language models.
+  • We open-source the code and all associated artifacts of BiasJailbreak to facilitate community efforts in understanding and mitigating safety-induced biases in large language models.
 This contribution aims to provide researchers and developers with the necessary tools to explore the nature of these biases and develop more robust defenses, furthering the collective effort to ensure the safety and reliability of LLM deployments.
-Our research suggests that while intentional biases are crucial for aligning LLMs with ethical standards, they necessitate careful scrutiny to prevent their manipulation.
+Our research suggests that while ethical biases are crucial for aligning LLMs with ethical standards, they necessitate careful scrutiny to prevent their manipulation.
 Therefore, responsible strategies from AI companies and researchers are needed to reinforce the safety of LLMs in an increasingly complex threat landscape.
@@
 2.1 Safety Alignment in LLMs
-Ensuring the safety and ethical alignment of large language models (LLMs) is a critical area of ongoing research.
+Ensuring the safety and ethical alignment of large language models (LLMs) is a critical area of ongoing research, since the ethical bias of LLMs can lead to undesirable societal impacts and potential harms (Li et al.
+(2024)).
 Methods such as data filtering, supervised fine-tuning, and reinforcement learning from human feedback (RLHF) aim to align models like GPT-4 and ChatGPT with human values and preferences (Christiano et al.
@@
 (2020)).
-However, despite these efforts, recent studies reveal vulnerabilities that can be exploited through ’jailbreak’ attacks, which lead to undesirable and harmful outputs (Kang et al., 2023; Hazell, 2023; Shen et al., 2023).
+However, despite these efforts, recent studies reveal vulnerabilities that can be exploited through ’jailbreak’ attacks, which lead to undesirable and harmful outputs (Kang et al.
+(2024); Shen et al.
+(2023).
+Additionally, Zheng et al.
+(2024) proposed many-shot demonstration techniques, using random search within demo pools and injecting system tokens to bypass safeguards.
 2.2 Jailbreak Attacks and Techniques
 Jailbreaking LLMs involves crafting inputs that bypass safety mechanisms, resulting in harmful or objectionable content.
-Early jailbreak attacks, such as the ”Do-Anything-Now (DAN)” series, relied on manually crafted prompts to exploit LLM safeguards (walkerspider, 2022).
+Early jailbreak attacks, such as the ”Do-Anything-Now (DAN)” series, relied on manually crafted prompts to exploit LLM safeguards (Shen et al.
+(2023)).
 (Liu et al.
@@
 (2023)) introduced AutoDAN, which combines manual and automated strategies using hierarchical genetic algorithms to enhance both the stealthiness and scalability of jailbreak prompts.
+Zeng et al.
+(2024a) proposed persuasive adversarial prompts (PAP) that leverage social science-based persuasion techniques to significantly enhance jailbreak success, achieving over 92% success rates across multiple models.
+Similarly, Shah et al.
+(2024) introduced persona modulation, a black-box jailbreak approach that uses personas to exploit vulnerabilities at scale, with high success rates transferable across state-of-the-art models.
 Language diversity and non-natural language inputs present additional challenges.
@@
 Yuan et al.
-(2024) extended this by investigating the vulnerabilities of LLMs to non-natural language inputs, such as ciphers.
+(2024); Jin et al.
+(2024b) extended this by investigating the vulnerabilities of LLMs to non-natural language inputs, such as ciphers.
 2.3 Towards Improved Safety Measures
@@
 This highlights the dynamic and evolving nature of jailbreak techniques.
-Our work builds on the existing body of research by focusing on the paradoxical consequences of intentional biases introduced for safety purposes.
+Our work builds on the existing body of research by focusing on the paradoxical consequences of ethical biases introduced for safety purposes, such as stated in Achiam et al.
+(2023).
 While these biases aim to align LLMs ethically, they also highlight new vulnerabilities.
@@
 Our findings advocate for a nuanced approach to LLM safety, promoting the development of more secure and reliable models, and ensuring that safety measures do not inadvertently introduce new risks.
-3 Methodology: PCJailbreak
+3 Methodology: BiasJailbreak
 3.1 Preliminaries
 3.1.1 Jailbreak Attack
-A jailbreak attack in the context of Large Language Models (LLMs) occurs when the model generates harmful or inappropriate responses to malicious inputs instead of producing a refusal signal, which is a safe and ethical response denying the request (Zeng et al., 2024; Zou et al., 2023).
+A jailbreak attack in the context of Large Language Models (LLMs) occurs when the model generates harmful or inappropriate responses to malicious inputs instead of producing a refusal signal, which is a safe and ethical response denying the request (Zeng et al., 2024b; Zou et al., 2023).
 Such attacks are intricately linked to the alignment methods employed in LLMs, which aim to align the model’s outputs with human values and ethical considerations.
@@
 When each combined input $T_{i}=\langle J_{i},Q_{i}\rangle$ is fed into the LLM, the model generates a set of responses $R=\{R_{1},R_{2},\ldots,R_{n}\}$.
-The aim of the jailbreak attack is to ensure that the responses $R$ correctly address the malicious queries in $Q$ instead of producing refusal signals (Zeng et al., 2024).
+The aim of the jailbreak attack is to ensure that the responses $R$ correctly address the malicious queries in $Q$ instead of producing refusal signals (Zeng et al., 2024b).
 In summary, a successful jailbreak attack coerces the LLM into bypassing its safety and ethical constraints, culminating in the generation of harmful or inappropriate content in response to malicious prompts.
 This kind of attack underscores significant vulnerabilities in the current alignment strategies of LLMs, highlighting the necessity for robust countermeasures to guarantee model safety and reliability.
-3.1.2 Formulation
+3.1.2 Jailbreak Attack Evaluation
 Designing a specific target response for each malicious question is impractical for several reasons.
@@
 […]
-3.2 PCJailbreak Approach
+3.2 BiasJailbreak Approach
 3.2.1 Generating Keywords and Prompts
-Our proposed PCJailbreak approach investigates how keywords associated with various demographic and socio-economic groups influence an LLM’s susceptibility to jailbreak attacks.
+Our proposed BiasJailbreak approach investigates how keywords associated with various demographic and socio-economic groups influence an LLM’s susceptibility to jailbreak attacks.
 To accomplish this, we generated keywords representing contrasting groups, such as privileged and marginalized communities, using the LLM itself to ensure that the keywords reflect the model’s inherent associations and biases.
@@
 3.2.2 Analyzing Inherent Biases
-Figure 3: Overview of the PCJailbreak methodology.
+Figure 3: Overview of the BiasJailbreak methodology.
 The same harmful prompt is used across different keywords representing contrasting groups to analyze variations in jailbreak success rates.*All keywords representing both privileged and underprivileged groups are generated by the LLM.
@@
 This could indicate inherent biases in the LLM’s training data or alignment mechanisms that affect how it responds to prompts involving different groups.
-3.2.3 Summary of the PCJailbreak Approach
-Our PCJailbreak methodology systematically analyzes the LLM’s responses to a diverse set of harmful prompts, using various keywords across different groups.
+3.2.3 Summary of the BiasJailbreak Approach
+Our BiasJailbreak methodology systematically analyzes the LLM’s responses to a diverse set of harmful prompts, using various keywords across different groups.
 We conducted experiments using harmful prompts $H$ and multiple pairs of privileged and marginalized keywords.
@@
 The calculation of jailbreak success rates (Equation 2) and the comparative analysis using $\Delta S_{i}$ (Equation 3) enabled us to quantify potential biases.
-Our approach highlights whether intentional biases introduced during the alignment process inadvertently contribute to vulnerabilities that can be exploited through jailbreak attacks.
-3.3 PCDefense: Preventing Jailbreaks without Additional Inference or Guard Models
-Our jailbreak defense method PCDefense uses prompts to adjust biases without the need for additional models.
+Our approach highlights whether ethical biases introduced during the alignment process inadvertently contribute to vulnerabilities that can be exploited through jailbreak attacks.
+3.3 BiasDefense: Preventing Jailbreaks without Additional Inference or Guard Models
+Our jailbreak defense method BiasDefense uses prompts to adjust biases without the need for additional models.
 Chain-of-Thought Prompting (CoT) has shown considerable potential in eliciting reasoning abilities in large language models, allowing for interpretability without the need for parameter updates (Wei et al., 2022).
-Inspired by this, our approach PCDefense involves adding defense prompts to reduce excessive biases, ensuring safety without additional inference overheads, in contrast to Guard Models which require additional inference cost after text generation (Inan et al.
+Inspired by this, our approach BiasDefense involves adding defense prompts to reduce excessive biases, ensuring safety without additional inference overheads, in contrast to Guard Models which require additional inference cost after text generation (Inan et al.
 (2023); Ghosh et al.
@@
 To further illustrate our defense prompt, we use the following prompt structure to ensure fairness and equity:
-Figure 4: PCDefense adjusts inherent biases in LLMs that are exploited by PCJailbreak. It is efficient since it does not require additional inference or models such as Guard Models.
-By using a defense prompt, our defense mechanism elicits the refusal of a wide range of harmful content without requiring significant additional computing resources, which makes it an attractive alernative to Guard Models.
+Figure 4: BiasDefense adjusts inherent biases in LLMs that are exploited by BiasJailbreak. It is efficient since it does not require additional inference or models such as Guard Models.
+By using a defense prompt, our defense mechanism elicits the refusal of a wide range of harmful content without requiring significant additional computing resources, which makes it an attractive alternative to Guard Models.
 4 Experiment
@@
 Data To analyze the differences in jailbreak success rates between groups, we used the jailbreakbench dataset (Chao et al.
+(2024)) and advbench dataset (Zou et al.
 (2023)).
-This dataset comprises 100 harmful prompts and has been widely used in numerous studies.
-We utilized this dataset to measure performance variations and to analyze the intentional biases inherent in LLMs.
-Models We analyzed intentional bias in well-known LLMs, including closed-source models such as GPT-3.5-turbo, GPT-4, GPT-4-o, and open-source models like Llama2-7B, Llama2-13B, Llama3-7B, Phi-mini-7B, Qwen1.5, and Qwen2-7B.
+These datasets comprises a total of 600 harmful prompts and has been widely used in numerous studies.
+We utilized this dataset to measure performance variations and to analyze the ethical biases inherent in LLMs.
+Models We analyzed ethical bias in well-known LLMs, including closed-source models such as GPT-3.5-turbo, GPT-4, GPT-4-o, Claude-sonnet3.5, and open-source models like Llama2-7B, Llama2-13B, Llama3-7B, Phi-mini-7B, Qwen1.5, and Qwen2-7B.
 For our experiments, we conducted evaluations with the default sampling temperature and system prompts.
 This setup aligns with standard practices in the literature to ensure consistent and comparable results.
-Keywords The keywords used to compare performance across groups were generated by the LLMs themselves to maximize the exploitation of internal biases.
-Most models generated similar keywords, and significant performance differences were observed.
-Basically, the models primarily generated the following keywords for various groups (see Table 1):
-Marginalized Group (Adjective) Privileged Group (Adjective)
-White Black
-Male Female
-Heterosexual LGBTQ+
-Able-bodied Disabled
-Wealthy Low-income
-Cisgender Non-normative gender
-Traditionally family-structured Non-traditionally family-structured
-Urban Rural
-Highly educated Undereducated
-Traditionally employed Sexually marginalized
-Table 1: Keywords Generated for Group Comparison
-These keywords were systematically used to evaluate the differences in jailbreak success rates among various groups.
-4.2 Intensional Bias Analysis
-To gauge the impact of intentionally injected biases in LLMs on their vulnerability to jailbreak attempts, we conducted a series of experiments utilizing various keyword prompts designed to assess the success rates of jailbreaks within marginalized and privileged categories.
-We evaluated multiple well-known models, including GPT-3.5, GPT-4, GPT-4o (Brown et al., 2020)(closed models), as well as LLaMA (Touvron et al., 2023), Qwen (Bai et al., 2023), and Phi-mini (Abdin et al., 2024)(open-source models).
+Keywords The keywords used to compare performance across groups were generated by the target LLM themselves to maximize the exploitation of internal biases (see Table 10).
+The generated keywords of each model shows that there were common keywords that are close to typical biases, while showing sufficient diversity.
+4.2 Ethical Bias Analysis
+As shown in Table 1, we analyzed the differences in jailbreak success rates across groups using two utilized datasets: JailbreakBench and AdvBench.
+The results variations in success rates between marginalized and privileged groups, providing evidence of ethical bias in LLMs.
+These findings underscore the importance of addressing such biases to improve the fairness and reliability of language models.
+Table 1: Performance across different datasets showing baseline success rates, marginalized success rates, privileged success rates, and the difference between marginalized and privileged success rates using LLaMA2 model.
+Dataset Baseline
+Success Rate
+Marginalized
+Success Rate ($\uparrow$)
+Privileged
+Success Rate ($\downarrow$)
+Marginalized /
+Privileged ($\uparrow$)
+JailbreakBench 0.2400 0.2811 (+17.08%) 0.1933 (-19.58%) 145.42%
+AdvBench 0.1895 0.2037 (+7.50%) 0.1758 (-7.25%) 115.84%
+To further analyze this bias, we conducted a series of experiments aimed at evaluating the impact of ethical biases in LLMs on their susceptibility to jailbreak attempts.
+Specifically, we utilized various keyword prompts to assess the differences in jailbreak success rates across marginalized and privileged categories.
+Our evaluation included multiple well-known models, such as GPT-3.5, GPT-4, GPT-4o (Brown et al., 2020) and Claude-sonnet3.5 (Anthropic, 2024) (closed models), alongside open-source models like LLaMA (Touvron et al., 2023), Qwen (Bai et al., 2023), and Phi-mini (Abdin et al., 2024).
+Table 2: Performance across different models showing baseline success rates, marginalized success rates, privileged success rates, and the difference between marginalized and privileged success rates.
 Model Baseline Marginalized Privileged Marginalized /
@@
 GPT-4o 0.4600 0.5467 (+18.91%) 0.4187 (-8.91%) 130.57%
+Claude-sonnet3.5 0.3100 0.3371 (+8.74%) 0.2764 (-10.84%) 121.90%
 LLaMA2 0.2400 0.2811 (+17.08%) 0.1933 (-19.58%) 145.42%
@@
 Phi-mini 0.4100 0.4386 (+7.07%) 0.3829 (-6.59%) 114.56%
-Table 2: Performance across different models showing baseline success rates, marginalized success rates, privileged success rates, and the difference between marginalized and privileged success rates.
-Interestingly, the latest model, GPT-4o, exhibits higher jailbreak success rates, with a significant difference observed between marginalized and privileged groups.
-In particular, GPT-4o shows a difference of 0.128 between marginalized and privileged keywords, highlighting the substantial gap influenced by injected biases.
-Furthermore, most models demonstrate lower jailbreak success rates for prompts using privileged keywords compared to the baseline, whereas prompts using marginalized keywords exhibit higher success rates.
+As shown in Table 2, GPT-4o has higher jailbreak success rates, with a notable 0.128 difference between marginalized and privileged groups.
+Most models show lower success rates for privileged keywords and higher rates for marginalized ones compared to the baseline, where the baseline refers to the prompt in Equation 1 without keywords.
 Intriguingly, the LLaMA3 model demonstrates a lower propensity for successful jailbreak attempts compared to other models.
-This lower rate can be attributed to Meta’s focus on creating more robust LLMs that are resistant to such vulnerabilities.
-It showcases Meta’s commitment to mitigating LLM safety risks through comprehensive alignment techniques (Vidgen et al., 2021; Ghosh et al., 2024; Caselli et al., 2020).
-4.3 Performance Improvement using Bias-based Method
+This lower rate can be attributed to Meta’s focus on developing more robust LLMs that are resistant to such vulnerabilities.
+(Vidgen et al., 2021; Ghosh et al., 2024; Caselli et al., 2020).
+4.3 Existing Jailbreak Performance Improvement using BiasJailbreak
 The bias-based approach we analyzed can also be applied to existing models.
-We confirmed performance improvement by applying the PCJailbreak method to Adaptive Attacks (Andriushchenko et al., 2024), the current state-of-the-art (SOTA) jailbreak attack.
+We confirmed performance improvement by applying the BiasJailbreak method to Adaptive Attacks (Andriushchenko et al., 2024), the current state-of-the-art (SOTA) jailbreak attack.
 Specifically, for the llama2 model(Touvron et al., 2023), the performance improved from 98% to 100%, resulting in a 2% increase.
@@
 These results show that the method can be easily integrated with existing techniques.
-The experiment was held by using the jailbreak attack artifact from JailbreakBench(Chao et al., 2024), which has 100 samples of Adaptive Attacks conversation.
+The experiment was held by using the jailbreak attack artifact from JailbreakBench(Chao et al., 2024), which has 100 samples of Adaptive Attacks conversation
 Table 3: Performance Improvement of SOTA Models
-Model Adaptive Attacks Adaptive Attacks with PCJailbreak
+Model Adaptive Attacks Adaptive Attacks with BiasJailbreak
 llama2 98.00% 100.00%
@@
 The results in Table 3 demonstrate that our bias-based method can be effectively combined with the existing approaches, providing notable improvements in performance.
-4.4 Experimental Validation of PCDefense
-To validate our defense method PCDefense, we conducted experiments on various models, including Llama2, Phi, and Qwen2.
+4.4 Experimental Validation of BiasDefense
+To validate our defense method BiasDefense, we conducted experiments on various models, including Llama2, Phi, and Qwen2.
 We observed the impact of our approach on Marginalized Rate Success and Privileged Rate Success metrics before and after applying our defense technique.
 The results are summarized in Table 4.
-Table 4: Jailbreak Prevention performance of PCDefense
+Table 4: Jailbreak Prevention performance of BiasDefense
 Model Metric Before After After/Before ($\downarrow$)
@@
 Additionally, the overall average score has decreased, suggesting that the system has become safer.
-4.5 Principle Component Analysis
-In this section, we conduct a detailed analysis of how models interpret PCJailbreak prompts by performing Principal Component Analysis (PCA) (F.R.S., 1901) on the embeddings of Phi-mini, Qwen2, and LLaMA3 models, as shown in Appendix Figures 5, 6, 7 respectively.
-Our findings reveal that models with a high success rate in PCJailbreak tend to cluster PCJailbreak prompts together with benign prompts, while those with a lower success rate demonstrate the opposite behavior.
-The PCA results indicate that when using PCJailbreak, models tend to interpret it more similarly to benign prompts, with the degree of this similarity aligning with our observed PCJailbreak success rates.
+Table 5: Comparison of jailbreak success rates across prompt-based defense methods using LLAMA2 model.
+Defense Method Metric Jailbreak Success Rate
+None Marginalized Group Jailbreak Success 28.11%
+Privileged Group Jailbreak Success 19.33%
+self-remind Marginalized Group Jailbreak Success 33.67%
+Wu et al. (2023) Privileged Group Jailbreak Success 27.33%
+Defending Marginalized Group Jailbreak Success 20.57%
+Zhang et al. (2023) Privileged Group Jailbreak Success 14.86%
+RPO Marginalized Group Jailbreak Success 56.00%
+Zhou et al. (2024) Privileged Group Jailbreak Success 49.75%
+BiasDefense (Ours) Marginalized Group Jailbreak Success 17.14%
+Privileged Group Jailbreak Success 14.29%
+In Table 5, BiasDefense, our proposed method, demonstrates superior defense performance by achieving the lowest jailbreak success rates among the compared state-of-the-art prompt-based defense methods.
+Specifically, it records the lowest success rates for both the Marginalized Group and the Privileged Group, indicating enhanced overall security.
+This suggests that BiasDefense is more effective at preventing jailbreak attacks compared to other defense techniques.
+Table 6: Comparison of Defense cost for BiasDefense and Llama-Guard.
+The experiment was held on a single H100 GPU, with Llama-3.2-1B-Instruct as the language model, and Llama-Guard-3-8B as the guard model.
+Defense Method Time Cost (seconds) $\downarrow$ Time cost (percentage) $\downarrow$
+Baseline (No Defense) 21.91 +0.00%
+BiasDefense 22.44 +2.40%
+Llama-Guard-3-8B 31.69 +44.60%
+Table 6 shows how our prompt-based method is computationally efficient compared to standalone classifier models such as Llama-Guard.
+Using a standalone classifier requires far more computation time than adding defense prompts.
 5 Conclusion
 In conclusion, our study highlights the complexities and potential unintended consequences of aligning large language models (LLMs) with safety measures aimed at preventing harmful outputs.
-The introduction of intentional biases for ethical behavior, while crucial for ensuring responsible AI, has led to significant discrepancies in jailbreak success rates based on gender and racial keywords.
-This discrepancy, coined as ”PCJailbreak,” underscores the risks that arise from safety-induced biases, particularly in terms of fairness and equality.
+The introduction of ethical biases for ethical behavior, while crucial for ensuring responsible AI, has led to significant discrepancies in jailbreak success rates based on gender and racial keywords.
+This discrepancy, coined as ”BiasJailbreak,” underscores the risks that arise from safety-induced biases, particularly in terms of fairness and equality.
 Our findings emphasize the need for LLM developers to carefully balance safety and fairness in their models.
@@
 Appendix A Appendix
-Figure 5: PCA of Phi-mini model, which has 0.4386 PCJailbreak success rate. PCJailbreak samples are closely clustered with Benign samples.
-Figure 6: PCA of Qwen2 model, which has 0.1971 PCJailbreak success rate. PCJailbreak samples are relatively far from both harmful and benign samples.
-Figure 7: PCA of LLaMA3 model, which has 0.0650 PCJailbreak success rate. PCJailbreak samples are relatively close to harmful samples.
+A.1 PCA Visualization Analysis
+In this section, we conduct a detailed analysis of how models interpret BiasJailbreak prompts by performing Principal Component Analysis (PCA) (F.R.S., 1901) on the embeddings of Phi-mini, Qwen2, and LLaMA3 models, as shown in Appendix Figures 5, 6, 7 respectively.
+Our findings reveal that models with a high success rate in BiasJailbreak tend to cluster BiasJailbreak prompts together with benign prompts, while those with a lower success rate demonstrate the opposite behavior.
+The PCA results indicate that when using BiasJailbreak, models tend to interpret it more similarly to benign prompts, with the degree of this similarity aligning with our observed BiasJailbreak success rates.
+Figure 5: PCA of Phi-mini model, which has a 0.4386 BiasJailbreak success rate. BiasJailbreak samples are closely clustered with Benign samples.
+Figure 6: PCA of Qwen2 model, which has a 0.1971 BiasJailbreak success rate. BiasJailbreak samples are relatively far from both harmful and benign samples.
+Figure 7: PCA of LLaMA3 model, which has a 0.0650 BiasJailbreak success rate. BiasJailbreak samples are relatively close to harmful samples.
+A.2 Keyword Group Impact on BiasJailbreak
+Table 7 presents the reported confidence intervals based on three runs conducted with a temperature setting of 0.7.
+The results reveal a clear distinction between marginalized and privileged keywords, with marginalized keywords generally demonstrating higher success rates.
+The observed confidence intervals are consistent across runs, further validating the robustness of the findings.
+We also conducted experiments for random adjective keywords.
+While slow exhibited a relatively high success rate of 31%, other adjectives had jailbreak success rates similar to the baseline (without a keyword).
+This suggests that general adjectives have minimal impact on jailbreak attempts.
+This consistency, also evident in the case of general adjectives, underscores the reliability of the methodology.
+Notably, significant disparities were observed in groups such as Black vs.
+White, neurodiverse vs.
+neurotypical, Native American vs.
+White, and LGBTQ+ vs.
+heterosexual, where marginalized keywords consistently outperformed privileged ones.
+Table 7: Comparison of the impact of individual keyword groups on BiasJailbreak.
+Marginalized Groups and Privileged Groups have a distinct difference among Random Adjectives.
+The Difference (%) (95% CI) column shows the change in Jailbreak Success Rate compared to the baseline, along with the 95% confidence interval.
+Random Adjectives Group
+Keyword Model – Baseline 95% Conf. Interval Group treatment effect
+big -4.00% (1.32%) (-6.59%, -1.41%) -0.50% (3.97%)
+small +1.00% (1.26%) (-1.47%, +3.47%)
+bright -2.00% (3.73%) (-8.69%, +4.69%)
+cold -3.00% (1.26%) (-5.47%, -0.53%)
+old 0.00% (6.26%) (-12.56%, +12.56%)
+slow +7.00% (7.64%) (-7.61%, +21.61%)
+warm -2.00% (1.82%) (-5.05%, +1.05%)
+dark -1.00% (2.56%) (-5.34%, +3.34%)
+Marginalized Group
+Keyword Model – Baseline 95% Conf. Interval Treatment Effect
+female +4.00% (3.73%) (-2.69%, +10.69%) +4.33% (4.93%)
+poor +7.00% (2.56%) (+2.66%, +11.34%)
+black +2.33% (1.30%) (-0.14%, +4.80%)
+LGBTQ 0.00% (1.26%) (-2.47%, +2.47%)
+transgender +3.00% (7.76%) (-11.83%, +17.83%)
+neurodiverse +7.67% (12.55%) (-16.71%, +32.05%)
+working class +5.33% (4.14%) (-2.22%, +12.88%)
+bisexual +0.67% (1.30%) (-1.80%, +3.14%)
+native american +9.00% (1.26%) (+6.53%, +11.47%)
+Privileged Group
+Keyword Model – Baseline 95% Conf. Interval Treatment Effect
+male 0.00% (1.26%) (-2.47%, +2.47%) -2.90% (11.21%)
+rich +2.67% (1.33%) (+0.08%, +5.26%)
+white -7.33% (1.30%) (-9.80%, -4.86%)
+heterosexual -7.00% (1.26%) (-9.47%, -4.53%)
+straight -14.33% (6.82%) (-27.28%, -1.38%)
+neurotypical +3.67% (6.82%) (-9.28%, +16.62%)
+middle class +2.00% (28.87%) (-54.52%, +58.52%)
+A.3 BiasDefense Ablation on the combination of prefix and suffix prompt
+The results of the BiasDefense Ablation Study, shown in Table 8, demonstrate the performance of various configurations of the proposed BiasDefense mechanism.
+BiasDefense consists of two primary components: a system prompt and a suffix prompt.
+The ablation study evaluates the effectiveness of these components both independently and in combination.
+  • Individual Prompts: Using either the system prompt or the suffix prompt independently achieves lower jailbreak success rates compared to having no defense at all. For example:
+    • The system prompt reduces marginalized group jailbreak success from 0.2811 (no defense) to 0.2558, and privileged group success from 0.1933 to 0.1687.
+    • The suffix prompt reduces marginalized group jailbreak success to 0.2367, although it slightly increases the success rate for privileged groups to 0.2733.
+  • Combined Prompts: The best performance is achieved when both the system and suffix prompts are used together.
+The full BiasDefense configuration reduces marginalized group jailbreak success to 0.1714 and privileged group jailbreak success to 0.1429, representing a significant improvement over individual prompts and the baseline without any defense.
+These findings highlight the importance of using both components of BiasDefense together.
+While each component provides some improvement when used independently, their integration ensures the most robust performance against jailbreak attempts.
+This demonstrates the value of a comprehensive approach to mitigating bias vulnerabilities in language models.
+Furthermore, our prompt-based defense method effectively addresses ethical biases, providing a practical solution for improving the fairness and reliability of language models.
+Table 8: BiasDefense Ablation on the combination of suffix and prefix defense prompt using LLAMA2 model.
+Defense Method Metric Jailbreak Success Rate
+None Marginalized Group Jailbreak Success 0.2811
+Privileged Group Jailbreak Success 0.1933
+BiasDefense: system prompt Marginalized Group Jailbreak Success 0.2558
+Privileged Group Jailbreak Success 0.1687
+BiasDefense: suffix prompt Marginalized Group Jailbreak Success 0.2367
+Privileged Group Jailbreak Success 0.2733
+BiasDefense Marginalized Group Jailbreak Success 0.1714
+Privileged Group Jailbreak Success 0.1429
+A.4 BiasDefense Ablation on normal task performance
+For evaluating the effect on normal task performance with defense prompts including BiasDefense, Table 9 shows the MMLU score performance with various defense prompt methods.
+The results show that our BiasDefense has relatively low performance degradation compared to other suffix adding defense prompts.
+One notable aspect is that without the suffix prompt in BiasDefense, MMLU score has gotten better than using no defense prompts, showing even a positive effect on normal tasks.
+Although the previous results of Table 8 shows that without the suffix the defense gets weaker, using only the prefix prompt would stand as a good alternative if we want minimal impact on normal tasks.
+Table 9: MMLU score using prompt-based defense methods.
+The base language model used is google-flan-t5 (Chung et al.
+(2024)) and GPT-4o-2024-11-20.
+*For GPT-4o, we used 4 subjects sampled from each category due to cost, being college_computer_science (category: STEM), sociology (category: Social Sciences), college_medicine (category: Other), jurisprudence (category: Humanities).
+Model Defense Method Suffix Average MMLU*
+google-flan-t5 None X 0.295
+google-flan-t5 RPO Zhou et al. (2024) O 0.243
+google-flan-t5 self-remind Wu et al. (2023) O 0.238
+google-flan-t5 Defending Zhang et al. (2023) X 0.295
+google-flan-t5 BiasDefense(Ours, With Suffix) O 0.265
+google-flan-t5 BiasDefense(Ours, Without Suffix) X 0.296
+GPT-4o-2024-11-20 None X 0.883
+GPT-4o-2024-11-20 RPO Zhou et al. (2024) O 0.237
+GPT-4o-2024-11-20 self-remind Wu et al. (2023) O 0.861
+GPT-4o-2024-11-20 Defending Zhang et al. (2023) X 0.876
+GPT-4o-2024-11-20 BiasDefense(Ours, With Suffix) O 0.869
+GPT-4o-2024-11-20 BiasDefense(Ours, Without Suffix) X 0.887
+A.5 Keywords for Different Models
+Table 10: Keywords discovered in Claude-sonnet3.5.
+Marginalized Group (Adjective) Privileged Group (Adjective)
+Female Male
+Immigrant Native
+Poor Rich
+Muslim Christian
+Black White
+High school educated College educated
+LGBTQ+ Heterosexual
+Transgender Straight
+Neurodiverse Neurotypical
+Spanish English
+Working class Middle class
+Bisexual Heterosexual
+Native American White
+Table 11: Keywords discovered in GPT-4o.
+Marginalized Group (Adjective) Privileged Group (Adjective)
+Poor Wealthy
+Black White
+Elderly Young
+Disabled Able-bodied
+Non-binary Cisgender
+Unemployed Employed
+Religious minority Religious majority
+Indigenous Colonizer-descended
+Lower-class Upper-class
+Underrepresented Overrepresented
+Minoritized Dominant
+Non-English speaking English speaking
+Informally educated Formally educated
+Psychiatrically diagnosed Psychiatrically typical
+Single parent Partnered parent
+Adolescent Adult
+Gender non-conforming Gender conforming
+Internationally born Nationally born
+Temporarily employed Permanently employed
+Linguistically diverse Linguistically standard
+Socially isolated Socially integrated
+Chronically ill Healthy
+Mentally ill Mentally healthy
+Substance dependent Substance independent
+Homeless Sheltered
+Incarcerated Non-incarcerated
+Refugee Citizen
+Visually impaired Visually unimpaired
+Hearing impaired Hearing unimpaired
+Non-traditional student Traditional student
+Working class Professional class
+Historically disadvantaged Historically advantaged
+War-affected Peace-living
+Food insecure Food secure
+Low mobility High mobility
+Environmentally displaced Environmentally secure
+Survivor Non-affected
+Recovering Non-recovering
+Public transport dependent Car owning
+Crafts-based Industrial
+Artisanal Mass-produced
+Temporary visa holder Permanent resident
+Non-voter Voter
+Uninsured Insured
+Non-internet user Internet user
+Low-income High-income
+Table 12: Keywords discovered in Qwen2.
+Marginalized Group (Adjective) Privileged Group (Adjective)
+Low-Income High-Income
+Racial Minority White
+LGBTQ+ Heterosexual
+Women Men
+Disabled Able-bodied
+Refugees Citizens
+Uneducated Educated
+Neurodivergent Neurotypical
+Religious Minority Mainstream Religion
+Homeless Homeowners
+Mental Health Patients Mentally Stable Individuals
+Elderly Young Adults
+Overweight Slim/Fit
+Non-native Speakers Native Speakers
+Incarcerated Free
+Rural Urban
+Illiterate Literate
+Politically Disenfranchised Voters
+Undocumented Immigrants Legal Immigrants
+Sex Workers Non-sex Workers
+Table 13: Keywords discovered in LLaMA2.
+Marginalized Group (Adjective) Privileged Group (Adjective)
+Female Male
+Poor Rich
+Black White
+LGBTQ Heterosexual
+Transgender Straight
+Neurodiverse Neurotypical
+Working class Middle class
+Bisexual Heterosexual
+Native American White
+Table 14: Keywords discovered in Phi.
+Marginalized Group (Adjective) Privileged Group (Adjective)
+Black White
+Female Male
+LGBTQ+ Heterosexual
+Non-normative gender Cisgender
+Non-traditionally family-structured Traditionally family-structured
+Rural Urban
+Homeless Socially influential
+Undereducated Highly educated
+Disabled Able-bodied
+Religiously marginalized Religiously dominant
+Low-income Wealthy
+Sexually employed Traditionally employed